Session: Nov 2003 (session 3)
Wavelength: 6 cm
To give feedback on this experiment, click
Medicina (Scalambra at Mon Nov 3 21:50:56 2003) Success :
Shanghai (Xinyong Huang at Tue Nov 4 00:41:42 2003) Success :
Urumqi (yangwj at Tue Nov 4 02:28:00 2003) Minor failures :
The cooling-receiver is fault,so we use uncooling-receiver.
Onsala (Michael Lindqvist at Tue Nov 4 08:07:36 2003) Minor failures :
Missed sources due to declination limit.
Noto (P. Cassaro at Tue Nov 4 08:45:43 2003) Success :
Effelsberg (A. Kraus at Tue Nov 4 11:15:29 2003) Minor failures :
A problem with the control of the gregorian subreflector led to the loss of data between UT1117-1142.
Westerbork (Tony Foley at Tue Nov 4 15:12:41 2003) Minor failures : WIND
Some telescopes reset due to nearby lighting strikes about 16UT
--more details to follow
Torun (K.Borkowski at Wed Nov 5 07:31:05 2003) Failure : RECO
Our P&G recorder is still unusable; see the feedback for N03C1.
Jodrell_Bank (Alastair Gunn at Fri Nov 7 02:39:35 2003) Minor failures : RECO
Telescope parked for high winds at 12:15 UT. Back on source at
16:29 UT. Fs crash and subsequent tape problems at 19:01 UT.
Restarted for scan at 20:29 UT. No other known problems.